# Tensor invariants of finite and classical groups

### v1

@@ Abstract.
Given a faithful representation $W$ of a finite group $G$, we classify when the invariant subring of the tensor algebra $T(W)$ is finitely generated/noetherian as a twisted commutative algebra (tca). For a representation $W$ of a classical group $G_{\mathbf{Z}}$, we show that the invariant subring $T(W_{k})^{G_{k}}$ is a finitely generated tca for $k$ algebraically closed of sufficiently large characteristics, assuming $W$ admits a good filtration over $\mathbf{Z}$. Finally, we introduce a categorical variant of the Gelfand–Kirillov dimension and compute its value to be $\binom{n+1}{2}$ for $T(\mathbf{C}^{n})$ as a tca.
@@ 1. Introduction
An overarching goal of this paper is to demonstrate how Schur–Weyl duality enables us to answer questions in noncommutative ring theory using standard methods from commutative algebra, even in positive characteristic. We specifically address the finite generation problem for tensor invariants by relating it to Weyl’s polarization technique using the Schur functor. Throughout this paper, we let $k$ be an infinite field of characteristic $p\geq 0$.
@@
Given a finite group $G$ acting on a finite-dimensional vector space $W$, it is well known that the invariant subring $\operatorname{Sym}(W)^{G}$ is finitely generated. The analogous question for the tensor algebra $T(W)$ is quite subtle with a rich history. Wolf [Wol36] showed that the ring of noncommutative symmetric polynomials in $\geq 2$ variables is not finitely generated. Almost half a century later, Dicks–Formanek [DF83] proved this almost always holds: the subring of tensor invariants $T(W)^{G}$ is a finitely generated $k$-algebra if and only if $G$ acts via a finite cyclic group consisting of scalar matrices.
Kharčenko recognized the significance of the commuting actions of the symmetric groups $\mathfrak{S}_{n}$ on $T(W)_{n}$ to this question. He [Kha84] asked whether $T(W)^{G}$ admits a finite generating set when incorporating this action, noting that this holds for soluble groups $G$. Koryukin [Kor84] answered Kharčenko’s question affirmatively for linearly reductive groups. To that end, he introduced the notion of an “S-algebra” and proved a noetherian result for $T(W)$ as an S-algebra. From this, he deduced that $T(W)^{G}$ is also a finitely generated S-algebra using the Reynolds operator. Over the past decade, Sam–Snowden [SS12, SS17] and others have independently studied S-algebras under the name “twisted commutative algebras” (tcas), and we adopt this terminology exclusively. See Section 2.2 for definitions.
@@
We settle all remaining cases of the finite generation problem for finite group tensor invariants in the negative while further strengthening Koryukin’s positive result.
@@
  • is noetherian and generated by elements of degree #$\leq\#G$, if $p\nmid\#G$ or $p=0$; and
  • is not generated in finitely many degrees, if $p\mid\#G$.
The infinite generation of the invariant ring for the permutation action of $\mathfrak{S}_{p}$ on $T(k^{p})$ as a tca when $p>0$ was obtained only recently in [BDDK23] (see Example 3.6). Our contribution to the first part is twofold: we establish a module-theoretic variant of Koryukin’s theorem where, in our terminology, he only shows that $T(W)^{G}$ is weakly noetherian; and we make Koryukin’s result effective.
Our key idea is to bridge the study of tensor invariants with the invariants of certain commutative—albeit infinite-dimensional—algebras via Schur–Weyl duality and then invoke results from classical invariant theory. We recall relevant results about the Schur functor in Section 2. In Section 3, we prove Theorem 1.1 by transporting results of Richman [Ric96] about the failure of Weyl’s polarization theorem for finite groups in the modular setting and of Fleischmann extending Noether’s bound to small characteristics [Fle00] to the noncommutative case.
While the first part of our main theorem without the generation bound extends easily to infinite linearly reductive groups, the second part does not immediately generalize to infinite reductive groups. We partially address this with our next result. Let $G$ be a classical group and $W$ be a finite $\mathbf{Z}$-module with a linear action of $G$ affording a universal good filtration; see Section 4.1. The base change $W_{k}=W\otimes k$ is a representation of $G_{k}=G\times_{\mathbf{Z}}k$. By carrying over results of Derksen–Makam [DM21] through our bridge, we obtain:
@@
In Example 4.2, we show that the conclusion of the above theorem can fail for small positive values of $p$. The structure of the invariant rings occurring in Theorem 1.2 is of great interest, as determining whether or not they are noetherian is a major open problem in the field of representation stability (see [Gan24] and references therein). We thus prove some auxiliary results about their structure (Theorem 4.3).
In the final section, we introduce the Gelfand–Kirillov dimension for finitely generated algebra objects in a locally finite length symmetric monoidal category and compute it for some tcas. Notably, in characteristic zero, $\operatorname{GKdim}$ of polynomial tcas generated in degree one agrees with the Krull–Gabriel dimension of their module category as computed by Nagpal–Sam–Snowden [SS19, Section 3.5]. It would be interesting to prove a general result along these lines.
@@ 1.3. Related work
Tambour and Teranishi independently pioneered the use of Schur–Weyl duality to study tensor invariants, albeit only in characteristic zero. In [Ter88], Teranishi explains how to explicitly obtain generators of $T(W)^{G}$ as a tca using generators of $\operatorname{Sym}(W^{\oplus\dim W})^{G}$, a method made possible by combining Schur–Weyl duality with Weyl’s polarization theorem. He [Ter91] also provides a novel method to derive generators of $T(W)^{G}$ from that of $\operatorname{Sym}(W\oplus{\textstyle\bigwedge}^{2}W)^{G}$. In his thesis [Tam92], Tambour reproves Koryukin’s theorem, obtains degree bounds for the invariant ring, and introduces and proves rationality results for a certain enhanced Hilbert series by using classical results from invariant theory like Noether’s bound and Molien’s formula.
Twisted (commutative) algebras, even infinitely generated ones, arise naturally when studying the topology of configuration spaces of points/discs on a plane (see for example, [AM24, Waw23] and references therein). We hope the method developed in Section 5 can be used to provide interesting invariants of topological spaces.
@@ 2. Setup
Throughout this paper, a partition $\lambda=(\lambda_{1},\lambda_{2},\ldots)$ is an infinite sequence of integers with $\lambda_{i}\geq\lambda_{i+1}$ for all $i\geq 1$ and $\lambda_{j}=0$ for $j\gg 0$. The size of the partition $|\lambda|=\sum_{i}\lambda_{i}$. A partition is $p$-restricted if $\lambda_{i}-\lambda_{i+1}<p$ for all $i\geq 1$. We also let $\mathbf{V}$ be the infinite-dimensional $k$-vector space with basis $e_{1},e_{2},\ldots$ and $\mathbf{GL}$ be the group of $k$-linear automorphisms of $\mathbf{V}$.
@@ 2.1. Symmetric monoidal categories
For definitions about symmetric monoidal categories and functors, we refer the reader to [Rei16, Chapter 2]. We first introduce the two symmetric monoidal categories that are central to this paper.
@@
    $\operatorname{Rep}^{\mathrm{pol}}(\mathbf{GL}):$ A representation of $\mathbf{GL}$ is polynomial if it occurs as a subquotient of direct sums of $\mathbf{V}^{\otimes n}$ with $n$ allowed to vary. We let $\operatorname{Rep}^{\mathrm{pol}}(\mathbf{GL})$ denote the category of polynomial representations of $\mathbf{GL}$; it is a symmetric monoidal category with the usual tensor product of representations (denoted $\otimes$). The simple objects in this category are indexed by arbitrary partitions $\lambda$ and denoted $L_{\lambda}(\mathbf{V})$.
@@
    $(V\boxtimes W)_{n}=\bigoplus_{i=0}^{n}\operatorname{Ind}_{\mathfrak{S}_{i}\times\mathfrak{S}_{n-i}}^{\mathfrak{S}_{n}}V_{i}\boxtimes W_{n-i}.$
    This endows $\operatorname{Rep}(\mathfrak{S}_{*})$ with the structure of a symmetric monoidal category. The simple objects in this category are indexed by $p$-restricted partitions $\lambda$ and denoted $D^{\lambda}$.
The above two categories have a canonical $\mathbf{N}$-grading with $L_{\lambda}(\mathbf{V})$ (and $D^{\lambda}$) homogeneous of degree $|\lambda|$. The monoidal structure preserves this grading on both categories.
@@ 2.2. Commutative algebra objects
A $\mathbf{GL}$-algebra is a commutative algebra object in $\operatorname{Rep}^{\mathrm{pol}}(\mathbf{GL})$. Concretely, a $\mathbf{GL}$-algebra $S$ is a commutative $k$-algebra $S$ with an action of $\mathbf{GL}$ by $k$-algebra automorphisms under which $S$ forms a polynomial representation of $\mathbf{GL}$. ∎
@@
A twisted commutative algebra (tca) is a commutative algebra object in $\operatorname{Rep}(\mathfrak{S}_{*})$. Concretely, a tca is an associative, unital, $\mathbf{N}$-graded $k$-algebra $A=\bigoplus_{n\geq 0}A_{n}$ equipped with an action of $\mathfrak{S}_{n}$ on $A_{n}$ such that:
@@
The last condition above is the “twisted commutativity” condition. ∎
Let $W$ be a $k$-vector space considered as a representation of $\mathfrak{S}_{1}$; we emphasize this by instead writing $W\langle 1\rangle$. The polynomial tca over $W\langle 1\rangle$ is the non-commutative polynomial ring $T(W)$ over $W$ with the $S_{n}$-action on $T(W)_{n}$ being the obvious one: on monomials it is given by $\sigma(w_{1}\cdots w_{n})=w_{{\sigma^{-1}(1)}}\cdots w_{{\sigma^{-1}(n)}}$.

### v2

# Noncommutative invariants of finite and classical groups

@@ Abstract.
[…]
@@ 1. Introduction
An overarching goal of this paper is to demonstrate how Schur–Weyl duality enables us to answer questions in noncommutative ring theory using standard methods from commutative algebra, even in positive characteristic. We specifically address the finite generation problem for invariants of tensor algebras by relating it to Weyl’s polarization technique using the Schur functor. Throughout this paper, we let $k$ be an infinite field of characteristic $p\geq 0$.
@@
Given a finite group $G$ acting on a finite-dimensional vector space $W$, it is well known that the invariant subring $\operatorname{Sym}(W)^{G}$ is finitely generated. The analogous question for the tensor algebra $T(W)$ is quite subtle with a rich history. Wolf [Wol36] showed that the ring of noncommutative symmetric polynomials in $\geq 2$ variables is not finitely generated. Almost half a century later, Dicks–Formanek [DF83] proved that this almost always holds: the subring of noncommutative invariants $T(W)^{G}$ is a finitely generated $k$-algebra if and only if $G$ acts via a finite cyclic group consisting of scalar matrices.
Kharčenko recognized the significance of the commuting actions of the symmetric groups $\mathfrak{S}_{n}$ on $T(W)_{n}$ to this question. He [Kha84] asked whether $T(W)^{G}$ admits a finite generating set when incorporating this action, noting that this holds for soluble groups $G$ when $p=0$. Koryukin [Kor84] answered Kharčenko’s question affirmatively for linearly reductive groups. To that end, he introduced the notion of an “S-algebra” and proved a noetherian result for $T(W)$ as an S-algebra. From this, he deduced that $T(W)^{G}$ is also a finitely generated S-algebra using the Reynolds operator. Over the past decade, Sam–Snowden [SS12, SS17] and others have independently studied S-algebras under the name “twisted commutative algebras” (tcas), and we adopt this terminology exclusively. See Section 2.3 for definitions.
@@
We negatively resolve the finite generation problem for noncommutative invariants of finite groups in the modular setting.
@@
  • is noetherian and generated by elements of degree $\leq\#G$, if $p\nmid\#G$ or $p=0$; and
  • is not finitely generated, if $p\mid\#G$.
The infinite generation of the invariant ring for the permutation action of $\mathfrak{S}_{n}$ on $T(k^{n})$ as a tca when $n\geq p>0$ was obtained only recently in [BDDK23] (see Example 3.6). Our contribution to the non-modular setting (the first part) is limited to the following: we establish a module-theoretic noetherian result, whereas, in our terminology, Koryukin only shows that $T(W)^{G}$ is weakly noetherian, and we make his result effective.
Our key idea is to bridge the study of noncommutative invariants with the invariants of certain commutative—albeit infinite-dimensional—algebras using Schur–Weyl duality and then invoke recent results from modular invariant theory. This setup is explained in detail in Section 2, but we sketch the idea here. Let $S=\operatorname{Sym}(W\otimes\mathbf{V})$ and $R=T(W)$ where $W$ is a representation of a group $G$ and $\mathbf{V}\coloneqq\operatorname{colim}k^{n}$ is an infinite-dimensional vector space over $k$. The objects $S$ and $R$ can be viewed as free commutative algebra objects in two important categories from representation theory, namely $\operatorname{Rep}^{\mathrm{pol}}(\mathbf{GL})$ and $\operatorname{Rep}(\mathfrak{S}_{*})$. Furthermore, $S$ is mapped to $R$ under the Schur functor $\mathcal{S}\colon\operatorname{Rep}^{\mathrm{pol}}(\mathbf{GL})\to\operatorname{Rep}(\mathfrak{S}_{*})$. Our key observation is:
Theorem 1.2.
Under the identification $\mathcal{S}(S)\cong R$ above,
  • we have an isomorphism $\mathcal{S}(S^{G})\cong R^{G}$, and
  • given a $\mathbf{GL}$-subrepresentation $U\subset S^{G}$ that generates $S^{G}$, the subspace $\mathcal{S}(U)\subset R^{G}$ generates $R^{G}$.
In Section 3, we first prove Theorem 1.2. The second part of Theorem 1.1 now follows by transporting results of Richman [Ric96] regarding the failure of Weyl’s polarization theorem for finite groups in the modular setting. Similarly, the first part of Theorem 1.1 follows by invoking Fleischmann’s work, which extends Noether’s bound to all characteristics [Fle00], combined with standard arguments involving the Reynolds operator.
While the first part of our main theorem extends easily to infinite linearly reductive groups, the second part does not immediately generalize to infinite reductive groups. We partially address this with our next result. Let $G$ be a classical group and $W$ be a finite $\mathbf{Z}$-module with a linear action of $G$ affording a universal good filtration; see Section 4.1. The base change $W_{k}=W\otimes k$ is a representation of $G_{k}=G\times_{\mathbf{Z}}k$. By carrying over results of Derksen–Makam [DM21] through the bridge from Theorem 1.2, we obtain:
@@
The conclusion of the above theorem can fail for small $p>0$ (Example 4.2).
In the final section, we introduce the Gelfand–Kirillov dimension for finitely generated algebra objects in a locally finite length symmetric monoidal $k$-linear category and compute it for some tcas. Notably, in characteristic zero, $\operatorname{GKdim}$ of polynomial tcas generated in degree one agrees with the Krull–Gabriel dimension of their module category as computed by Nagpal–Sam–Snowden [SS19, Section 3.5]. It would be interesting to prove a general result along these lines.
@@ 1.3. Related work
Tambour and Teranishi independently pioneered the use of Schur–Weyl duality to study noncommutative invariants, albeit only in characteristic zero. In [Ter88], Teranishi explains how to explicitly obtain generators of $T(W)^{G}$ as a tca using generators of $\operatorname{Sym}(W^{\oplus\dim W})^{G}$, a method made possible by combining Schur–Weyl duality with Weyl’s polarization theorem. He [Ter91] also provides a novel method to derive generators of $T(W)^{G}$ from those of $\operatorname{Sym}(W\oplus{\textstyle\bigwedge}^{2}W)^{G}$. In his thesis [Tam92], Tambour reproves Koryukin’s theorem, obtains degree bounds for the invariant ring, and introduces and proves rationality results for a certain enhanced Hilbert series by using classical results from invariant theory like Noether’s bound and Molien’s formula.
The structure of the invariant rings occurring in Theorem 1.3 is of great interest, as determining whether or not they are noetherian is a major open problem in the field of representation stability (see [Gan24, Gan25] for related results). While we do not address this in the present paper, we prove some auxiliary results (Theorem 4.3).
Twisted (commutative) algebras, even infinitely generated ones, arise naturally when studying the topology of configuration spaces of points/disks on a plane (see, for example, [AM24, Waw23] and references therein). We hope the method developed in Section 5 can be used to provide interesting invariants of topological spaces.
Acknowledgments.
The author thanks Mátyás Domokos, Nate Harman, Martin Kassabov, and Andrew Snowden for helpful discussions over email and is especially grateful to Lucas Buzaglo for patiently explaining foundational results about noncommutative rings.
@@ 2. Setup
Throughout this paper, a partition $\lambda=(\lambda_{1},\lambda_{2},\ldots)$ is an infinite sequence of integers with $\lambda_{i}\geq\lambda_{i+1}$ for all $i\geq 1$ and $\lambda_{j}=0$ for $j\gg 0$. The size of the partition $|\lambda|=\sum_{i}\lambda_{i}$. A partition is $p$-restricted if $\lambda_{i}-\lambda_{i+1}<p$ for all $i\geq 1$. We also let $\mathbf{V}$ be the infinite-dimensional $k$-vector space with basis $e_{1},e_{2},\ldots,$ and $\mathbf{GL}$ be the group of $k$-linear automorphisms of $\mathbf{V}$.
@@ 2.1. Symmetric monoidal categories
We first sketch the definition of a symmetric monoidal $k$-linear category and functors between them. A symmetric monoidal $k$-linear abelian category is the data $(M,\otimes,1,\alpha,\lambda,B)$ where:
@@
    $M$ is a $k$-linear abelian category
  2. (2)
    $\otimes$ is a $k$-bilinear functor
    $\otimes:M\times M\to M$
    called the tensor product;
  3. (3)
    $1\in M$ is an object called the unit object; and
  4. (4)
    we have natural isomorphisms
    $a_{x,y,z}$ $:(x\otimes y)\otimes z\to x\otimes(y\otimes z)$ (associator)
    $\lambda_{x}$ $:1\otimes x\to x$ (left unitor)
    $B_{x,y}$ $:x\otimes y\to y\otimes x$ (braiding)
such that $B_{x,y}B_{y,x}=\mathrm{id}_{y\otimes x}$ and certain coherence axioms are satisfied (the associator obeys the pentagonal identity; the associator and unitor obey the triangle identity; and the braiding and associator obey the hexagonal identity). For the full coherence axioms, we refer the reader to [Rei16, Chapter 1.3].
Given two symmetric monoidal $k$-linear abelian categories $(M,\otimes_{M},1_{M},a^{M},\lambda^{M},B^{M})$ and $(N,\otimes_{N},1_{N},a^{N},\lambda^{N},B^{N})$, a symmetric monoidal $k$-linear functor $M\to N$ is the data $(F,\varphi,\epsilon)$ where
  1. (1)
    $F\colon M\to N$ is a $k$-linear functor;
  2. (2)
    $\varphi$ is a natural isomorphism
    $\varphi_{x,y}:F(x)\otimes_{N}F(y)\xrightarrow{\sim}F(x\otimes_{M}y),\quad\forall x,y\in M,$
  3. (3)
    $\epsilon$ is an isomorphism $1_{N}\xrightarrow{\sim}F(1_{M})$,
and these morphisms are also required to satisfy certain compatibility conditions (with the associativity, unitor, and braiding morphisms).
For the remainder of the paper, we will suppress the natural maps and the tensor product when referring to a symmetric monoidal $k$-linear category. Given a $k$-linear functor $F$ between two symmetric monoidal $k$-linear categories, we say it is symmetric monoidal if there exists an isomorphism $\epsilon$ and a natural isomorphism $\varphi$ such that $(F,\varphi,\epsilon)$ is a symmetric monoidal $k$-linear functor as defined above.
2.2. Representation categories
We introduce the two symmetric monoidal $k$-linear categories that are central to this paper.
  1. (1)
    $\operatorname{Rep}^{\mathrm{pol}}(\mathbf{GL}):$ A representation of $\mathbf{GL}$ is polynomial if it occurs as a subquotient of direct sums of $\mathbf{V}^{\otimes n}$, with $n$ allowed to vary. We let $\operatorname{Rep}^{\mathrm{pol}}(\mathbf{GL})$ denote the category of polynomial representations of $\mathbf{GL}$; it is a symmetric monoidal category with the usual tensor product of representations (denoted $\otimes$) and the usual braiding on vector spaces.
    The simple objects in $\operatorname{Rep}^{\mathrm{pol}}(\mathbf{GL})$ are indexed by arbitrary partitions $\lambda$ (corresponding to their highest weight) and denoted $L_{\lambda}(\mathbf{V})$. Of particular interest are also the Schur modules, denoted $\mathbf{S}_{\lambda}(\mathbf{V})$; the simple module $L_{\lambda}(\mathbf{V})$ is the socle of $\mathbf{S}_{\lambda}(\mathbf{V})$.
@@ 2.2. Representation categories
    $(V\boxtimes W)_{n}=\bigoplus_{i=0}^{n}\operatorname{Ind}_{\mathfrak{S}_{i}\times\mathfrak{S}_{n-i}}^{\mathfrak{S}_{n}}V_{i}\otimes W_{n-i},$
    where $V_{i}\otimes W_{n-i}$ is the external tensor product of the $\mathfrak{S}_{i}$-representation and the $\mathfrak{S}_{n-i}$-representation nin-i. This endows $\operatorname{Rep}(\mathfrak{S}_{*})$ with the structure of a symmetric monoidal category where the braiding is the one induced from the braiding on vector spaces.
    The simple objects in $\operatorname{Rep}(\mathfrak{S}_{*})$ are indexed by $p$-restricted partitions $\lambda$ and denoted $D_{\lambda}$. We will also be interested in the Specht modules $S^{\lambda}$ which are the base changes of the integrally defined Specht modules via $\mathbf{Z}\to k$. Similar to the $\mathbf{GL}$-case, we have $D_{\lambda}=\operatorname{soc}(S^{\lambda})$.
We caution that the standard indexing of irreducible representations of the symmetric groups is by “$p$-regular partitions” and denoted $D^{\lambda}$; this is conjugate to our indexing, and in fact, $D^{\lambda}=D_{\lambda^{t}}\otimes\operatorname{sgn}$, where $\operatorname{sgn}$ is the sign representation of $\mathfrak{S}_{|\lambda|}$. This subtlety is not relevant to the contents of our paper; the interested reader may refer to [Gre07, §6.4].
Both categories have an $\mathbf{N}$-grading where $L_{\lambda}(\mathbf{V})$ and $D_{\lambda}$ are homogeneous of degree $|\lambda|$; the monoidal structure is compatible with this grading.
@@ 2.3. Commutative algebra objects
Given a symmetric monoidal $k$-linear category $\mathcal{C}$, a commutative algebra object $(A,\mu,\eta)$ is an object $A$ in $\mathcal{C}$ with a unit map $\eta\colon 1_{\mathcal{C}}\to A$ and a multiplication map $\mu\colon A\otimes A\to A$ such that the following diagrams commute (corresponding respectively to the associativity, unit, and commutativity axioms):
$(A\otimes A)\otimes A$$A\otimes(A\otimes A)$$A\otimes A$$A\scriptstyle{a_{A,A,A}}$$\scriptstyle{\mu\otimes 1_{A}}$$\scriptstyle{1_{A}\otimes\mu}$$\scriptstyle{\mu}$$1\otimes A$$A{A\otimes A}$$\scriptstyle{\lambda_{A}}$$\scriptstyle{\eta\otimes 1_{A}}$$\scriptstyle{\mu}$$A\otimes A$$A\otimes A$$A\scriptstyle{B_{A,A}}$$\scriptstyle{\mu}$$\scriptstyle{\mu}$
We will, by slight abuse of notation, refer to a commutative algebra object simply by its underlying object. Similar to the above discussion, given a commutative algebra object $A$, one can also define the notion of a module object for $A$ in $\mathcal{C}$.
We are now prepared to define our main objects of study.
@@
A $\mathbf{GL}$-algebra is a commutative algebra object in $\operatorname{Rep}^{\mathrm{pol}}(\mathbf{GL})$. ∎
Concretely, a $\mathbf{GL}$-algebra $S$ is an (ordinary) commutative $k$-algebra $S$ with an action of $\mathbf{GL}$ by $k$-algebra automorphisms under which $S$ forms a polynomial representation of $\mathbf{GL}$.
@@
A twisted commutative algebra (tca) is a commutative algebra object in $\operatorname{Rep}(\mathfrak{S}_{*})$. ∎
Concretely, a tca is an associative, unital, $\mathbf{N}$-graded $k$-algebra $A=\bigoplus_{n\geq 0}A_{n}$ equipped with an action of $\mathfrak{S}_{n}$ on $A_{n}$ such that:
@@
The last condition above is the “twisted commutativity” condition.
Let $W$ be a $k$-vector space considered as a representation of $\mathfrak{S}_{1}$; we emphasize this by instead writing $W\langle 1\rangle$. The free tca over $W\langle 1\rangle$ is the non-commutative polynomial ring $T(W)$ over $W$ with the $S_{n}$-action on $T(W)_{n}$ being the obvious one: on monomials, it is given by $\sigma(w_{1}\cdots w_{n})=w_{{\sigma^{-1}(1)}}\cdots w_{{\sigma^{-1}(n)}}$.